# Trapping of Single Atoms in Metasurface Optical Tweezer Arrays

### v2

@@ Methods
From the ultracold strontium gas, we typically load optical tweezer arrays at a trap depth of 100 µK. The trapping light is provided by a 520 nm 5W fiber laser with second harmonic generation (Azurlight, ALS-GR-520-5-A-CP-SF), seeded by a home-built extended cavity diode laser operating at 1040 nm (QPhotonics, QLD-1030-100S). Directly after the 520 nm laser output, we use an acousto-optic modulator to dynamically control the trap depth of the tweezers. In front of the metasurface, we use a magnifying telescope to increase the beam waist to be larger than the area of the metasurface. After the metasurface, a high-power capable microscope lens ($\mathrm{NA}=0.6$, Thorlabs, LMH-50X-532) is used to collimate the generated pattern. The tweezer pattern is relayed through a 1:1 telescope before being focused down onto the atoms via an objective lens ($\mathrm{NA}=0.5$, Mitutoyo, G Plan Apo 50X).
[…]
@@
Extended Data Fig. 4: Real image of a metasurface and the beam profile after diffraction. a, Real size comparison between an American one-cent coin and a 2.32 mm $\times$ 2.32 mm metasurface. b, The profile of a 1.5 mm diameter beam that is incident onto the metasurface. c, Far-field imaging of non-diffracted light after aligning the metasurface to the beam. This diagnostic, in tandem with optimizing the diffracted power, ensures our system is well aligned.
@@
To determine how well we can distinguish one from zero atoms in a tweezer trap, we use a model-free approach that does not require detailed modeling of the photon counting statistics. Our methodology is similar to the one used in Ref. [49], involving the recording of two fluorescence images of the atom array in short sequence, separated by 25 ms, within the same iteration of the experiment. Besides the detection fidelity, $F$, it yields the initial filling fraction $f$ and survival rate $S$ of atoms between images.
From the two images, the atomic fluorescence signal (photon count) at each trap location is recorded. For each trap, there are four possible outcomes and associated probabilities: there is an atom in both images ($p_{11}$), there is an atom in the first but not in the second image ($p_{10}$), there is an atom in the second but not first image ($p_{10}$), and there is no atom present in either image ($p_{00}$). The probability for each event is given by:
@@
where $F_{0}$ denotes the detection fidelity for zero atoms, $F_{1}$ the detection fidelity of one atom, $S$ is the survival rate between images, and $f$ the initial filling fraction. Initially, the threshold photon count $x$ that marks the distinction between zero and one atom is a variable. By iteratively varying $x$ (see Extended Data Fig. 5), we solve the above equations to maximize the total imaging fidelity defined by $F=fF_{1}+(1-f)F_{0}$. By optimizing the threshold $x$ for each individual tweezer, we find a mean detection fidelity across the array of $F=99.8\%$, a mean survival rate of $84.1\%$, and a mean filling fraction of $49.2\%$. The reduced survival rate $S$ is attributed to an enhanced loss process associated with the 520 nm trapping wavelength (see main text).
Extended Data Fig. 5: Determination of the photon count threshold $x$ to distinguish between zero and one atom in a trap. Data points show the counts for one specific trap of the $4\times 4$ array in 500 iterations of the experiment. A darker color of the data points indicates a higher density of points. The threshold value $x$ divides the data into four quadrants, labeled by $p_{ij}$ with indices $i=0,1$ and $j=0,1$ indicating the absence (presence) of an atom in the first and second image, respectively.
Extended Data Fig. 6: Measurements of uniformity in the $16\times 16$ array. a, Measurement of the trap depth via the light shift $\Delta$ of the ${}^{1}S_{0}$ ${}^{3}P_{1}$ resonance in the presence of the trapping field. Data shows the resonance for a single trap. b, Positional displacement between observed and target trap locations. c-d, show the measurement of the radial and axial trap frequency, respectively, via parametric heating. All error bars in this figure show $1\sigma$ s.e.m. from 20 repetitions of the measurement.
@@
The trap depth is measured via the light shift on the trapped atoms induced by the trapping laser. We probe the light shift on the $m_{j}=\pm 1$ transition between the 1S0 and 3P1 states using the 689 nm MOT beams. When the probe light is resonant with the shifted resonance, the atoms are heated out of the trap. Extended Data Fig. 6a shows a typical loss resonance in an individual trap. Fitting the resonance feature with a Gaussian model, we determine the detuning $\Delta$ from the free space resonance. From this we obtain the trap depth $U=\Delta\alpha_{{}^{1}\mathrm{S}_{0}}/(\alpha_{{}^{3}\mathrm{P}_{1}}-\alpha_{{}^{1}\mathrm{S}_{0}})$, where $\alpha_{{}^{1}\mathrm{S}_{0}}$ ($\alpha_{{}^{3}\mathrm{P}_{1}}$) is the polarization of the 1S0 (3P1) state.
The positional accuracy of the tweezer array is extracted by comparing the measured trap locations with the target trap positions that were used for the design of the metasurface. The trap locations are obtained by fitting each atom’s fluorescence signal to a 2D Gaussian. The orientation of the metasurface is not fixed to the lab frame and can rotate freely. We minimize the averaged distance between the measured and target trap locations by rotating and translating the coordinate pattern. Following this optimization, we find that the mean deviation of trap locations is 60 nm. As shown in Extended Data Fig. 6b the deviations between measured and target locations are uniformly distributed in all directions, indicating that there is no systematic bias.
The trap frequencies are measured via parametric heating. Individual trap loss features at twice the radial and axial trap frequencies are shown in Extended Data Figs. 6c and d, respectively. The parametric resonance in the radial direction shows a slight asymmetry. There are two distinct trap frequencies in the radial direction, which we attribute to the trapping potential not being perfectly circularly symmetric. We fit the line shape with a double Gaussian and plot the geometric mean as the radial trap frequency. The parametric resonance in the axial direction shows a symmetric dip, which we fit with a single Gaussian. Extended Data Fig. 7 shows the axial trap frequencies in each trap and their statistical spread.
@@
Extended Data Fig. 8: Calculation of the effective NA of a pixel-based beam shaping device. a, The phase profile of a lens with focal length $f$ is emulated by a pixelated phase mask. The individual pixels have a size $d$. A flat wavefront impinges on the device and it is converted into a focusing wavefront. $\theta_{\mathrm{m}}$ is the maximal angle for which the phase advance between neighboring pixels stays smaller than $\pi/2$. The gray scale indicates the phase shift of the metasurface pixels. b, Zooming into the wavefront advance of individual pixels, following Huygens’ principle, the angle $\theta_{\mathrm{m}}$ is determined by the pixel where the wavefront advance between neighboring pixels reaches $\lambda/4$ (corresponding to a phase shift of $\pi/2$). $\Delta\theta$ denotes the angular separation between neighboring pixels. The color scale indicates the phase shift of the emerging wavefront behind the metasurface.
@@
For our metasurface, $d/\lambda\sim 0.6$ and the effective NA calculated within this simplified model is 0.41, while the actual NA for our metasurfaces is 0.6. Thanks to the high NA, holographic metasurfaces can directly generate diffraction-limited traps on the micrometer scale in their image plane. For current liquid crystal SLMs, the effective NA is 0.05 or less due to their relatively large pixel size (4 µm – 20 µm). They require high quality demagnification optics with a high NA to realize tweezer arrays on the micrometer scale. Such optics can act as an unintended low-pass filter, cutting off high-frequency spatial Fourier components, posing a challenge to generate uniform and well-resolved tweezer arrays. While such issues can be partially mitigated by iterative optimization of the SLM pattern, holographic metasurfaces offer a way to fundamentally circumvent such issues.
@@
To numerically simulate the holographic performance of pixel-based beam shaping devices, we employ a Fast Fourier Transform (FFT)-based numerical method to solve the Rayleigh-Sommerfeld diffraction integral for a given phase-only hologram with specified pixel size and resolution. This method has been highly effective in designing optimal phase-only holograms and we have found excellent agreement with our experimental results. We analyze holograms with varying pixel sizes, starting with the smallest pixel size of a commercially available liquid crystal SLM (3.74 µm, GAEA-2 SLM from HOLOEYE Photonics AG) and extending down to the subwavelength pixel size of the metasurfaces used in this work (290 nm). For each pixel size, the hologram is set to a fixed resolution of $300\times 300$ pixels. The position of the focal plane is calculated based on the hologram’s physical size ($300d\times 300d$) and the effective NA. Each hologram is optimized using the modified Gerchberg-Saxton algorithm to generate a $3\times 3$ square array of traps with 5 µm spacing in the focal plane without intermediate optics. The optimized hologram serves as the input for the final propagation step in our simulation. By averaging the full width at half maximum (FWHM) values along the $x$ and $y$ axes (assuming the optical axis is along $z$ axis) of the optical intensity profile for each trap, we extract the average and range of FWHM values across the entire array, as presented in Fig. 5c. Our results demonstrate that subwavelength pixel sizes enable the direct generation of diffraction-limited traps with FWHMs smaller than the working wavelength.
@@
Furthermore, we numerically investigate the achievable uniformity of trap intensities while increasing the array size. We calculate the normalized standard deviation of trap intensities as a function of the number of generated tweezers in the focal plane for device resolutions ranging from $1000\times 1000$ to $000\times 16,$ 000 × 16 , 000 pixels. The simulation assumes a pixel size of $d=290$ nm and an $\mathrm{NA}=0.6$, corresponding to the metasurfaces used in this work. The position of the focal plane is determined by the hologram size (resolution $\times$ pixel size) and the NA. The holograms are optimized using the algorithm outlined above, and a final forward propagation was performed using the optimized holograms to simulate the tweezer array formed in the focal plane. We integrate the intensity profile for each tweezer and calculate the standard deviation for the entire array based on these values.
As shown in Fig. 5d, we find that for each resolution, the uniformity starts to decay rapidly beyond a certain number of tweezers. From the data we find empirically that at least 300 pixels per tweezer trap are needed to maintain a uniformity $>95\%$.

### v3

@@ Methods
From the ultracold strontium gas, we typically load optical tweezer arrays at a trap depth of 100 µK. The trapping light is provided by a 5-W, 520-nm fiber laser with second harmonic generation (Azurlight, ALS-GR-520-5-A-CP-SF), seeded by a home-built extended cavity diode laser operating at 1040 nm (QPhotonics, QLD-1030-100S). Directly after the 520-nm laser output, we use an acousto-optic modulator to dynamically control the trap depth of the tweezers. In front of the metasurface, we use a magnifying telescope to increase the beam waist to be larger than the area of the metasurface. After the metasurface, a high-power-capable microscope lens ($\mathrm{NA}=0.6$, Thorlabs, LMH-50X-532) is used to collimate the generated pattern. The tweezer pattern is relayed through a 1:1 telescope before being focused down onto the atoms via an objective lens ($\mathrm{NA}=0.5$, Mitutoyo, G Plan Apo 50X).
We perform fluorescence imaging to detect the atomic occupation of the tweezer array. We resonantly scatter photons on the 461-nm transition for 50 ms, collect the fluorescence through the high-NA objective that focuses the tweezers, separate the light from the tweezer path via a long-pass dichroic, and image on an EMCCD camera (Andor iXon Ultra 888). We image with a 200-mm lens before the camera, such that a single camera pixel corresponds to a real space size of 260 nm×260 $\text{ nm}\times 260\text${ nm}.
@@
Extended Data Fig. 4: Images of a metasurface, a laser beam profile, and the beam profile after diffraction by the metasurface. a, A photo of a 2.32 mm $\times$ 2.32 mm metasurface by the side of an American one-cent coin. b, The profile of a 1.5-$m$ diameter beam that is incident onto the metasurface. c, A far-field image of non-diffracted light after aligning the metasurface to the beam. This diagnostic, in tandem with maximizing the diffracted power, ensures good optical alignment.
Extended Data Fig. 5: Single atom preparation and detection in a $\mathbf{4\times 4}$ metasurface tweezer array. a, Average of 100 fluorescence images after parity projection. The high uniformity indicates that all traps have an approximately equal chance of being filled with an atom. b, Individual fluorescence image. Trap locations are indicated by dashed boxes. c, Histogram of the number of occupied sites after parity projection; the mean occupancy of a trap, marked by the dashed line, is $49(3)\%$. d, Determination of the photon count threshold $x$ to distinguish between zero and one atom in a trap. Data points show the photon counts for one specific trap in 500 iterations of the experiment. A darker color indicates a higher density of points. The threshold value $x$ divides the data into four quadrants, labeled by $p_{ij}$ with indices $i=0,1$ and $j=0,1$ indicating the absence (presence) of an atom in the first and second image, respectively. e, Histogram of photon counts across the trap locations of the $4\times 4$ array, as marked in b. 500 repetitions of the experiment are averaged. The data allows for the distinction between sites with one and zero atoms with high fidelity. The dashed line marks the averaged threshold value across the array.
@@
To determine how well we can distinguish one from zero atoms in a tweezer trap, we use a model-free approach that does not require detailed modeling of the photon counting statistics. Here, we discuss the approach for a $4\times 4$ array, with the corresponding data shown in Extended Data Fig. 5. For the $16\times 16$ array in the main text, we use the same methodology. The larger array shows a somewhat reduced single-atom detection fidelity compared to the smaller array which is due to limited available laser power rather than imperfections of the metasurface hologram. Empirically, we find that $\sim 7.5$ mW of power per tweezer trap provide optimal imaging conditions, while only $\sim 4$ mW are available for the $16\times 16$ array.
Our methodology is similar to the one used in Ref. [72], involving the recording of two fluorescence images of the atom array in short sequence, separated by 25 ms, within the same iteration of the experiment. Besides the detection fidelity, $F$, it yields the initial filling fraction $f$ and survival rate $S$ of atoms between images. From the two images, the atomic fluorescence signal (photon count) at each trap location is recorded. For each trap, there are four possible outcomes and associated probabilities: there is an atom in both images ($p_{11}$), there is an atom in the first but not in the second image ($p_{10}$), there is an atom in the second but not first image ($p_{10}$), and there is no atom present in either image ($p_{00}$). The probability for each event is given by:
@@
where $F_{0}$ denotes the detection fidelity for zero atoms, $F_{1}$ is the detection fidelity of one atom, $S$ is the survival rate between images, and $f$ is the initial filling fraction. Initially, the threshold photon count $x$ that marks the distinction between zero and one atom is a variable. By iteratively varying $x$ (see Extended Data Fig. 5d), we solve the above equations to maximize the total imaging fidelity defined by $F=fF_{1}+(1-f)F_{0}$. By optimizing the threshold $x$ for each individual tweezer, we find a mean detection fidelity across the array of $F=99.8\%$, a mean survival rate of $84.1\%$, and a mean filling fraction of $49.2\%$. The reduced survival rate $S$ is attributed to an enhanced loss process associated with the 520-nm trapping wavelength (see main text).
@@
The trap depth is measured via the light shift on the trapped atoms induced by the trapping laser. We probe the light shift on the $m_{j}=\pm 1$ transition between the 1S0 and 3P1 states using the 689-nm MOT beams. When the probe light is resonant with the shifted resonance, the atoms are heated out of the trap. Extended Data Fig. 6a shows a typical loss resonance in an individual trap. Fitting the resonance feature with a Gaussian model, we determine the detuning $\Delta$ from the free space resonance. From this we obtain the trap depth $U=\Delta\alpha_{{}^{1}\mathrm{S}_{0}}/(\alpha_{{}^{3}\mathrm{P}_{1}}-\alpha_{{}^{1}\mathrm{S}_{0}})$, where $\alpha_{{}^{1}\mathrm{S}_{0}}$ ($\alpha_{{}^{3}\mathrm{P}_{1}}$) is the polarization of the 1S0 (3P1) state. We report the statistical variance rather than the measurement uncertainty as it is the main contribution to the total uncertainty.
The positional accuracy of the tweezer array is extracted by comparing the measured trap locations with the target trap positions that were used for the design of the metasurface. The trap locations are obtained by fitting each atom’s fluorescence signal to a 2D Gaussian distribution. The positional inaccuracy between intended and measured locations is defined as $\Delta r=\sqrt{\Delta x^{2}+\Delta y^{2}}$, where $\Delta x$ ($\Delta y$) is the deviation in the $x$ ($y$) direction. The orientation of the metasurface is not fixed to the lab frame and can rotate freely. We minimize the averaged distance between the measured and target trap locations by rotating and translating the coordinate pattern. Following this optimization, we find that the mean deviation of trap locations is 60 nm. As shown in Extended Data Fig. 6b the deviations between measured and target locations are uniformly distributed in all directions, indicating that there is no systematic bias. The deviation is also comparable to the extent of the ground-state vibrational wavefunctions of Sr of about 30 nm and 90 nm in the radial and axial direction, respectively.
Extended Data Fig. 6: Measurements of uniformity in the $\mathbf{16\times 16}$ array. a, Measurement of the trap depth via the light shift $\Delta$ of the $^{1}S_{0} -$ $^{3}P_{1}$ resonance in the presence of the trapping field. Data shows the resonance for a single trap. b, Positional displacement between observed and target trap locations. c-d, show the measurement of the radial and axial trap frequency, respectively, via parametric heating. All error bars in this figure show $1\sigma$ s.e.m. from 20 repetitions of the measurement.
The trap frequencies are measured via parametric heating. The intensity of the tweezer light is modulated sinusoidally with an amplitude of $5\%$ for 30 ms while the frequency of the modulation is varied. Individual trap loss features at twice the radial and axial trap frequencies are shown in Extended Data Figs. 6c and d, respectively. The parametric resonance in the radial direction shows a slight asymmetry. There are two distinct trap frequencies in the radial direction, which we attribute to the trapping potential not being perfectly circularly symmetric. We fit the line shape with a double Gaussian distribution and plot the geometric mean as the radial trap frequency. The parametric resonance in the axial direction shows a symmetric dip, which we fit with a single Gaussian distribution. Extended Data Fig. 7 shows the axial trap frequencies in each trap and their statistical spread. We report the statistical variance rather than the measurement uncertainty as it is the main contribution to the total uncertainty.
@@
Extended Data Fig. 8: Calculation of the effective NA of a pixel-based beam shaping device. a, The phase profile of a lens with focal length $f$ is emulated by a pixelated phase mask. The individual pixels have a size $d$. A flat wavefront impinges on the device and it is converted into a focusing wavefront. $\theta_{\mathrm{m}}$ is the maximal angle for which the phase advance between neighboring pixels stays smaller than $\pi/2$. The gray scale indicates the phase shift of the metasurface pixels. b, Zooming into the wavefront advance of individual pixels, following Huygens’ principle, the angle $\theta_{\mathrm{m}}$ is determined by the condition where the wavefront advance between neighboring pixels reaches $\lambda/4$ (corresponding to a phase shift of $\pi/2$). $\Delta\theta$ denotes the angular separation between neighboring pixels. The color scale indicates the phase shift of the emerging wavefront behind the metasurface.
@@
For our metasurface, $d/\lambda\sim 0.6$ and the effective NA calculated using this simplified model is 0.41, while the actual NA for our metasurfaces is 0.6. Thanks to the high NA, holographic metasurfaces can directly generate diffraction-limited traps on the micrometer scale in their image plane. For current liquid-crystal SLMs, the effective NA is 0.05 or less due to their relatively large pixel size (4 µm – 20 µm). They require high-quality demagnification optics with a high NA to realize tweezer arrays on the micrometer scale. Such optics can act as an unintended low-pass filter, cutting off high-frequency spatial Fourier components and thus posing a challenge to generate uniform and well-resolved tweezer arrays. While such issues can be partially mitigated by iterative optimization of the SLM pattern, holographic metasurfaces offer a way to fundamentally circumvent such issues.
Extended Data Fig. 9: Imaging the $\mathbf{600\times 600}$ tweezer array. a, Composite image of the full array, stitched together from 126 individual high-resolution images. Each image captures a real-space area of $190\times 140$ µm. The individual images show a discontinuity in the measured light intensity, which arises from imperfections in the imaging system and contributes to the reported tweezer non-uniformity. b, Averaged tweezer spots from the center and the edges of the array. Each picture averages $\sim 300$ traps. Tweezers at the array edges show a pinching as well as a weak halo oriented towards the center of the array.
@@
To numerically simulate the holographic performance of pixel-based beam shaping devices, we employ a fast Fourier transform (FFT)-based numerical method to solve the Rayleigh-Sommerfeld diffraction integral for a given phase-only hologram with specified pixel size and resolution. This method is highly effective in designing and optimizing phase-only holograms, and yields excellent agreement between calculations our experimental results. We analyze holograms with varying pixel sizes, starting with the pixel size of a commercially available liquid-crystal SLM (3.74 µm, GAEA-2 SLM from HOLOEYE Photonics AG) and decreasing to the subwavelength pixel size of the metasurfaces used in this work (290 nm). For each pixel size, the hologram is set to a fixed resolution of $300\times 300$ pixels. The position of the focal plane is calculated based on the hologram’s physical size ($300d\times 300d$) and the effective NA. Each hologram is optimized using the modified Gerchberg-Saxton algorithm to generate a $3\times 3$ square array of traps with a 5-µm spacing in the focal plane without intermediate optics. By averaging the full width at half maximum (FWHM) values along the $x$ and $y$ axes (assuming the optical axis is along $z$ axis) of the optical intensity profile for each trap, we extract the average and range of FWHM values across the entire array, as presented in Fig. 5c. Our results demonstrate that subwavelength pixel sizes enable the direct generation of diffraction-limited traps with FWHMs smaller than the working wavelength.
@@
Furthermore, we numerically investigate the achievable uniformity of trap intensities while increasing the array size. We calculate the normalized standard deviation of trap intensities as a function of the number of generated tweezers in the focal plane for device resolutions ranging from $1000\times 1000$ to $16,000\times 16,000$ pixels. The simulation assumes a pixel size of $d=290$ nm and an $\mathrm{NA}=0.6$, corresponding to the metasurfaces used in this work. The position of the focal plane is determined by the hologram size (resolution $\times$ pixel size) and the NA. The holograms are optimized using the algorithm outlined above, and a final forward propagation is performed using the optimized holograms to calculate the tweezer array formed in the focal plane. We integrate the intensity profile for each tweezer and calculate the standard deviation for the entire array.
As shown in Fig. 5d, we find that for each resolution, the uniformity starts to decay rapidly beyond a certain number of tweezers. From the data we find empirically that at least 300 pixels per tweezer trap are needed to maintain a uniformity $95\%$.
$\mathbf{600\times 600}$ tweezer array characterization
The $600\times 600$ array spans an area of 1.5 mm2 in the image plane. Capturing a single image of the entire array while resolving individual tweezers with 2.5 µm spacing is challenging. We perform a raster scan and section the entire array into 126 high-resolution images that are recorded using a high-NA microscope ($\text{NA}=0.85$). These images are then stitched together to create a composite image of the full array. In order to reduce edge effects of the high-NA microscope, each image overlaps with neighboring ones by several rows and columns. A high-resolution composite image of the entire $600\times 600$ array and a video zooming into and out of the array are provided in the Supplementary Information.
As discussed in the main text, we measure an intensity non-uniformity across the array of $8\%$. This is an upper bound for the actual non-uniformity of the array. As can be seen in Fig. 9a, individual images that contribute to the full image consistently show an intensity gradient that is introduced by the imaging optics rather than the metasurface itself. We do not correct for such imaging-induced imperfections; they are included in the calculation of non-uniformity. To assess the roundness of individual tweezers at different locations in the large array, we show averaged tweezer spots from the array center and the edges, see Fig. 9b. While tweezers at the edges of the array show a weak halo and some pinching, pointing away from the center of the array, tweezers near the center show a high degree of roundness.